OpenStudy (anonymous):

The width of a rectangular computer screen is 5 inches more than its height. If the area of the screen is 234 square inches, find its dimensions.

OpenStudy (anonymous):

take width as w and height as y.so w=5+y.now area =width *height.so u get the equation y^2+5*y-234=0.solve it .u will get height y.find w from y.